OpenStudy (anonymous):

find the are and circumference of a circle with diameter 8yd

OpenStudy (anonymous):

c = pi*d c = 3.14 * 8 a = pi * r^2 a = 3.14 * 4^2

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

I'm doing that but my online math course is still saying it's incorrect and I need to use 3.14 which is \[\pi\]

OpenStudy (precal):

your c=251.2 a=50.24 read the instructions, what place value do they want for the decimal solutions.